OpenStudy (anonymous):

ok .. this integral was in my test today ... !!! x^3ln(1-x^(-4)) ... how would you solve it ? because i didn't =[

OpenStudy (anonymous):

your teacher must hate you

OpenStudy (mr.math):

Use integration by parts.

OpenStudy (anonymous):

this is a set up for integration by parts, but really i think it will take several steps

OpenStudy (anonymous):

i managed to get rid of the x^3 in the test .. but the i was stuck with the integral ln(1-1/x)

OpenStudy (anonymous):

oh maybe not \[u=\ln(1-x^{-4}), du = \frac{-4}{x-x^5}, dv = x^3, v = \frac{x^4}{4}\] maybe it is not so bad

OpenStudy (anonymous):

second integral turns in to \[\int\frac{x^3}{x^4-1}dx\] so a u -sub cures taht one

OpenStudy (anonymous):

maybe we will see the chart method.

OpenStudy (amistre64):

i was stuck on \[\int 2x^3 e^{2x}dx\]:) i couldnt parse it correctly on the test to save my life

OpenStudy (amistre64):

e^(x^2) not e^(2x)

OpenStudy (anonymous):

what is the chart method ? and how did you get to that integral ? can you show ?

OpenStudy (anonymous):

amistre makes a chart, i am not sure how

OpenStudy (amistre64):

i use crayons lol

OpenStudy (anonymous):

i was pretty sure you were not allowed sharp objects...

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

second integral is \[\int v du\] so it is \[\frac{x^4}{4}\times \frac{-4}{x-x^5}dx=-\int \frac{x^3}{1-x^4}dx\]

OpenStudy (anonymous):

i think the problem may have been you were thinking of reducing the power on x^3, when instead you should focus on getting rid of the log

OpenStudy (amistre64):

since the product rule is the basis for integration by parts; can it be adapted for multiple products? \[\int fgh = something profound?\]

OpenStudy (anonymous):

good question. will come back to it